# Detectability of dark matter subhalo impacts in Milky Way stellar streams

### v2

@@ 3.1.5 Summary
The discussion in this section gives us the ingredients to calculate total errors for the 5 observables ($\rho/\rho_{0}$, $\overline{\Delta\theta}$, $\overline{\Delta v_{r}}$, $\overline{\Delta v_{t}}$, and $\overline{\Delta v_{z}}$) we use for subhalo impacts. The total errors are given by Eq. 16 for $\rho/\rho_{0}$, Eq. 17 for $\overline{\Delta\theta}$, and Eq. 14 for $\overline{\Delta v_{r}}$, $\overline{\Delta v_{t}}$ and $\overline{\Delta v_{z}}$, with the effective uncertainty $\sigma_{X}^{\star 2}$ given by Eq. 21. These equations in turn use the number of stars per bin $n$ given by Eq. 15, and the fraction of observable stars $n_{\mathrm{obs}}/n$ under different magnitude cutoffs as summarized in Fig. 3.
The effective uncertainty $\sigma_{X}^{\star 2}$ depends also on stream width $\sigma_{\theta}$ and the stream distance $r_{0}$. Some results for $\sigma_{X}^{\star 2}$ are already shown in the bottom panels of Fig. 4 and Fig. 5. In order to perform tests of detectability in all of the stream parameter space, we calculate a denser grid of $\sigma_{X}^{\star 2}$. The results are shown in Fig. 6 for different observables and surveys, with an overall trend of smaller errors for colder and nearby streams. The effective errors on proper motions for Gaia are similar to Gaia+LSST (see the bottom left panel of Fig. 5) and thus not shown. Also, the effective errors for $v_{t}$ are similar to $v_{z}$ and thus not shown.
@@ 4 Detectability of DM subhalos
 $M_{\rm sh}$ Subhalo mass $[0,10^{9}]~{}\mathrm{M}_{\odot}$
@@
 $\psi_{\mathrm{impact}}$ Impact location $0^{\circ}$ on a $[-22^{\circ},22^{\circ}]$ region $[-4^{\circ},4^{\circ}]$
@@
where $\nu_{ij}$ denotes the value of the $i$th observable for the $j$th $\psi$ bin along the stream. We use a bin size of 2 degrees by default. The data consists of a mock impact $\nu_{ij}^{\rm dat}$ as described by Sec. 2.3, along with the total error in that bin $\sigma_{\nu_{ij}}$ as summarized in Sec. 3.1.5. Note that the error $\sigma_{\nu_{ij}}$ is computed assuming the number of stars per bin $n$ is constant across bins, so that the error $\sigma_{\nu_{ij}}$ is also constant across bins.
@@
Without this assumption, there would be an exact degeneracy between subhalo mass $M_{\mathrm{sh}}$ and relative velocity $\bm{w}_{\mathrm{rel}}=(w_{r},w_{\parallel},w_{z})$ in the analytic model, where taking $M_{\mathrm{sh}}\to\lambda M_{\mathrm{sh}}$ and $\bm{w}_{\mathrm{rel}}\to\lambda\bm{w}_{\mathrm{rel}}$ leaves the velocity kicks Eqs. 1-3 unchanged (see also section 5.1 of Erkal & Belokurov (2015b)). Taking $r_{s}$ to be a function of $M_{\mathrm{sh}}$ breaks the degeneracy, and is essential for us to maximize the likelihood function over its parameters. In addition, we also introduce a new nuisance parameter $\psi_{\mathrm{impact}}$, which allows the impact location of the model to vary along the stream arm. Using Eq. 24, we can then define test statistics for an impact to be distinguishable from a smooth stream, as well as confidence intervals on $M_{\rm sh}$ for a detected impact.
Since the detectability of a subhalo impact depends on the nuisance parameters, we will set default values for these parameters when considering mock data. These are given in Tab. 3. We set an impact parameter $b=0$ corresponding to a direct impact as this is the best possible case, although the results do not vary significantly for $b\lesssim 200$ pc. We set the subhalo velocity $w=\sqrt{w_{r}^{2}+w_{t}^{2}+w_{z}^{2}}=180$ km/s as this is a typical DM velocity in the halo, and set the direction to be in the $\hat{z}$ direction. Note that here we are setting the DM velocity in the halo frame so that we can assign a typical value. The relative velocity in the frame of the stream, $w_{\mathrm{rel}}=\sqrt{w_{r}^{2}+(w_{t}-V_{c})^{2}+w_{z}^{2}}$, depends on the DM direction relative to the stream. For time $t$, we pick a default value that is representative of detectability for most values of $t$. The value of $t$ is scaled with stream distance $r_{0}$ in order for the observables at different $r_{0}$ to have the same orbital phase, yielding more comparable perturbation profiles. Lastly, $\psi_{\mathrm{impact}}=0^{\circ}$ is defined to be at the center of the impact in the data, and we consider a region of $44\degree$ ($[-22^{\circ},22^{\circ}]$) about the impact center in the likelihood. We find that this angular range is large enough to include the full signal for any impact from subhalos of masses up to $10^{9}\mathrm{M}_{\odot}$. In Sec. 4.4, we will explore the effect on detectability when we shrink the region considered, thus revealing the dependence on stream length. Variations with other default parameter choices will also be explored later in Sec. 4.5.
@@ 4.1 Minimum detectable subhalo mass
$q_{0}=\begin{cases}2\ln\frac{L(\hat{M}_{\mathrm{sh}},~{}\hat{\bm{\theta}})}{L(0)}&\hat{M}_{\mathrm{sh}}\geq 0\\ 0&\hat{M}_{\mathrm{sh}}<0\\ \end{cases}$ (26)
@@
We can then compute the test statistic $q_{0}$ for mock data generated using a true value of the subhalo mass $M_{\mathrm{sh}}^{\mathrm{true}}$, the default values of nuisance parameters mentioned before, and errors as summarized in Sec. 3.1.5. Fig. 7 shows the distribution of $q_{0}$ as a function of $M_{\mathrm{sh}}^{\mathrm{true}}$, for a stream similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$, $\lambda=100~{}\mathrm{deg}^{-1}$) and different observational scenarios. The solid line is the median $q_{0}$ value for each $M_{\mathrm{sh}}^{\mathrm{true}}$, while the shaded area is the ±$\pm 1\sigma$ variation. These are obtained from 500 realizations of mock data at each $M_{\mathrm{sh}}^{\mathrm{true}}$. For each dataset, we evaluate $q_{0}$ using scipy.optimize.minimize to maximize the likelihood function over $M_{\mathrm{sh}}$ and $\bm{\theta}$. In particular, we use the Nelder-Mead (Gao & Han, 2012) method, with large boundaries on the value of each parameter, as specified in Tab. 3. Plots of the $q_{0}$ distribution at specific $M_{\mathrm{sh}}^{\mathrm{true}}$ values, as well as further discussion of the distribution shape can be found in App. A.
Figure 7: Test statistic $q_{0}$ for subhalo detectability (Eq. 26) as a function of true subhalo mass $M_{\mathrm{sh}}^{\mathrm{true}}$. Here we consider a stream similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$, $\lambda=100~{}\mathrm{deg}^{-1}$) for all four observational scenarios. The solid line indicates the median value of $q_{0}$ for each truth value of the subhalo mass, $M_{\mathrm{sh}}^{\mathrm{true}}$, and the band indicates the ±$\pm 1\sigma$ region. The dashed line indicates the threshold to reject the null hypothesis of no impact at 95% CL.
@@
In order to obtain the minimum detectable subhalo mass at 95% confidence level, we use the probability distribution for $q_{0}$ when the data $\nu_{ij}^{\mathrm{dat}}$ is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=0$ (no impact). The threshold for detectability is given by the 95th percentile value $q_{0}^{95\%}|_{M_{\mathrm{sh}}^{\mathrm{true}}=0}$, shown as the dashed lines in Fig. 7. For each observational scenario, the intersection of $q_{0}^{95\%}|_{M_{\mathrm{sh}}^{\mathrm{true}}=0}$ with the distribution of $q_{0}$ values for nonzero $M_{\mathrm{sh}}^{\mathrm{true}}$ gives the median and ±$\pm 1\sigma$ band for the minimum detectable subhalo mass at 95% CL.
Repeating this procedure for different streams, we can obtain the dependence of subhalo detectability on stream properties. Fig. 8 shows the minimum detectable subhalo mass as functions of stream width $\sigma_{\theta}$ (left), stream distance $r_{0}$ (middle) and stream density $\lambda$ (right) respectively. In each panel, we vary one stream property and keep the other two fixed. The fixed properties always take the values similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$ and $\lambda=$100 deg-1). The points in the plots correspond to the median while the error bars correspond to ±$\pm 1\sigma$. It can be seen that the minimum detectable mass increases as the stream width $\sigma_{\theta}$ increases, as the distance to the Galactic Center $r_{0}$ increases, or as the stream density $\lambda$ decreases. In addition, improvements in observational scenarios can significantly improve detectability.
@@
Figure 9: Decomposition of the test statistic $q_{0}$ into contributions from different observables. We assume a cold stream with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~{}\mathrm{kpc}$, $\lambda=100~{}\mathrm{deg}^{-1}$ and $q_{0}$ is evaluated on Asimov data sets, with no errors added, to show a representative breakdown. Top: We assume the minimum detectable subhalo mass in each observational scenario. Bottom: High subhalo mass $M_{\mathrm{sh}}=10^{8}~{}\mathrm{M}_{\odot}$ for all scenarios.
@@ 4.2 Fitting function for minimum detectable subhalo mass
$\log_{10}\frac{M_{\mathrm{sh}}^{\mathrm{min}}}{\mathrm{M}_{\odot}}=a\log_{10}\frac{\sigma_{\theta}}{\mathrm{deg}}+b\log_{10}\frac{r_{0}}{\mathrm{kpc}}+c\log_{10}\frac{\lambda}{\mathrm{deg}^{-1}}+d.$ (27)
@@
$a$ $b$ $c$ $d$
Gaia $1.19^{\pm 4\%}$ $1.91^{\pm 3\%}$ $-0.82^{\pm 4\%}$ $7.39^{\pm 1\%}$
DESI+Gaia $1.16^{\pm 4\%}$ $1.9^{\pm 3\%}$ $-0.84^{\pm 4\%}$ $7.42^{\pm 1\%}$
Via+Gaia+LSST $1.07^{\pm 5\%}$ $1.33^{\pm 5\%}$ $-0.83^{\pm 5\%}$ $7.31^{\pm 1\%}$
Via+Gaia+LSST10 $0.98^{\pm 9\%}$ $1.0^{\pm 12\%}$ $-0.81^{\pm 8\%}$ $7.32^{\pm 2\%}$
@@
From Tab. 4, we see that the detectability depends strongly on the stream properties for all observational scenarios. For example, an approximate scaling is given by $M_{\mathrm{sh}}^{\mathrm{min}}\propto\sigma_{\theta}^{1.2}r_{0}^{1.9}\lambda^{-0.8}$ for the Gaia scenario and $M_{\mathrm{sh}}^{\mathrm{min}}\propto\sigma_{\theta}^{0.98}r_{0}^{1.0}\lambda^{-0.8}$ for LSST 10 year sensitivity. In particular, the coefficient for the stream distance $r_{0}$, $b$, has very different values for the observational scenarios. The detectability under the first two scenarios degrades significantly ($b\approx 2$) with the increase of $r_{0}$, while in the latter two scenarios, the detectability degrades much more mildly ($b\approx 1$) with the increase of $r_{0}$. This is because the number of observable stars in Gaia decreases much faster with $r_{0}$ than for LSST 10 year sensitivity (see Fig. 3). The other coefficients have little or no dependence on observational scenario.
The application of this fit to a specific observed stream may not be that accurate because of the assumptions made in this work, including a) the circular orbit for the stream, b) the default parameters for the subhalo impact, and c) the assumption of fixed age and metallicity used in the MIST isochrones. However it gives us an effective way to quickly identify promising streams with potential for subhalo detection, from a large number of observed streams. We can always perform more accurate tests on the smaller set of promising streams that are selected from this approximate method.

### v3

@@ 3.1.5 Summary
The discussion in this section gives us the ingredients to calculate total errors for the 5 observables ($\rho/\rho_{0}$, $\overline{\Delta\theta}$, $\overline{\Delta v_{r}}$, $\overline{\Delta v_{t}}$, and $\overline{\Delta v_{z}}$) we use for subhalo impacts. The total errors are given by Eq. 14 for $\rho/\rho_{0}$, Eq. 15 for $\overline{\Delta\theta}$, and Eq. 21 for $\overline{\Delta v_{r}}$, $\overline{\Delta v_{t}}$ and $\overline{\Delta v_{z}}$, with the survey-effective uncertainty $\sigma_{X}^{\star 2}$ given by Eq. 20. These equations in turn use the number of stars per bin $n=\lambda\cdot(\psi{\rm binsize})$, and the fraction of observable stars $n_{\mathrm{obs}}/n$ under different magnitude cutoffs as summarized in Fig. 3.
The survey-effective uncertainty $\sigma_{X}^{\star}$ depends also on stream width $\sigma_{\theta}$ and the stream distance $r_{0}$. Some results for $\sigma_{X}^{\star}$ are already shown in the bottom panels of Fig. 4 and Fig. 5. In order to perform tests of detectability in all of the stream parameter space, we calculate a denser grid of $\sigma_{X}^{\star}$. The results are shown in Fig. 6 for different observables and surveys, with an overall trend of smaller errors for colder and nearby streams. The effective errors on proper motions for Gaia are similar to Gaia+LSST (see the bottom left panel of Fig. 5) and thus not shown. Also, the effective errors for $v_{t}$ are similar to $v_{z}$ and thus not shown.
@@ 4 Detectability of DM subhalos
$M_{\rm sh}$ Subhalo mass $0,10^{9}]~\mathrm{M}_{\odot}$
@@
$\psi_{\mathrm{impact}}$ Impact location $0^{\circ}$ on a $22^{\circ},22^{\circ}]$ region $4^{\circ},4^{\circ}]$
@@
where $\nu_{ij}$ denotes the value of the $i$th observable for the $j$th $\psi$ bin along the stream. We use a bin size of 2 degrees by default. We generate the mock impact data, $\nu_{ij}^{\rm dat}$, as described in Sec. 2.3, along with the total error in each bin $\sigma_{\nu_{ij}}$ as summarized in Sec. 3.1.5. Note that the error $\sigma_{\nu_{ij}}$ is computed assuming the number of stars per bin $n$ is constant across bins, so that the error $\sigma_{\nu_{ij}}$ is also constant across bins.
@@
Without this assumption, there would be an exact degeneracy between subhalo mass $M_{\mathrm{sh}}$ and relative velocity $\bm{w}_{\mathrm{rel}}=(w_{r},w_{\parallel},w_{z})$ in the analytic model, where taking $M_{\mathrm{sh}}\to\lambda M_{\mathrm{sh}}$ and $\bm{w}_{\mathrm{rel}}\to\lambda\bm{w}_{\mathrm{rel}}$ leaves the velocity kicks Eqs. 1-3 unchanged (see also section 5.1 of Erkal and Belokurov (2015b)). Taking $r_{s}$ to be a function of $M_{\mathrm{sh}}$ breaks the degeneracy, and is essential for us to maximize the likelihood function over its parameters. In addition, we also introduce a new nuisance parameter $\psi_{\mathrm{impact}}$, which allows the impact location of the model to vary along the stream arm. Using Eq. 24, we can then define test statistics for an impact to be distinguishable from a smooth stream, as well as confidence intervals on $M_{\rm sh}$ for a detected impact.
Since the detectability of a subhalo impact depends on the nuisance parameters, we will set default values for these parameters when considering mock data. These are given in Tab. 3. We set an impact parameter $b=0$ corresponding to a direct impact as this is the best possible case, although the results do not vary significantly when $b$ is smaller or at similar scale as the scale radius $r_{s}$. We set the subhalo velocity $w=\sqrt{w_{r}^{2}+w_{t}^{2}+w_{z}^{2}}=180$ km/s as this is a typical DM velocity in the halo, and set the direction to be in the $\hat{z}$ direction. Note that here we are setting the DM velocity in the halo frame so that we can assign a typical value. The relative velocity in the frame of the stream, $w_{\mathrm{rel}}=\sqrt{w_{r}^{2}+(w_{t}-V_{c})^{2}+w_{z}^{2}}$, depends on the DM direction relative to the stream. For time $t$, we pick a default value that is representative of detectability for most values of $t$. The value of $t$ is scaled with stream distance $r_{0}$ in order for the observables at different $r_{0}$ to have the same orbital phase, yielding more comparable perturbation profiles. Lastly, $\psi_{\mathrm{impact}}=0^{\circ}$ is defined to be at the center of the impact in the data, and we consider a region of $44\degree$ ($22^{\circ},22^{\circ}]$) about the impact center in the likelihood. We find that this angular range is large enough to include the full signal for any impact from subhalos of masses up to $10^{9}\mathrm{M}_{\odot}$. In Sec. 4.4, we will explore the effect on detectability when we shrink the region considered, thus revealing the dependence on stream length. Variations with other default parameter choices will also be explored later in Sec. 4.5.
@@ 4.1 Minimum detectable subhalo mass
[…]
@@
We can then compute the test statistic $q_{0}$ for mock data generated using a true value of the subhalo mass $M_{\mathrm{sh}}^{\mathrm{true}}$, the default values of nuisance parameters mentioned before, and errors as summarized in Sec. 3.1.5. Fig. 7 shows the distribution of $q_{0}$ as a function of $M_{\mathrm{sh}}^{\mathrm{true}}$, for a stream similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$, $\lambda=100~\mathrm{deg}^{-1}$) and different observational scenarios. The solid line is the median $q_{0}$ value for each $M_{\mathrm{sh}}^{\mathrm{true}}$, while the shaded area is the $\pm 1\sigma$ variation. These are obtained from 500 realizations of mock data at each $M_{\mathrm{sh}}^{\mathrm{true}}$. For each dataset, we evaluate $q_{0}$ by maximizing the likelihood function over $M_{\mathrm{sh}}$ and $\bm{\theta}$. In particular, we use the Nelder-Mead (Gao and Han, 2012) method implemented in scipy.optimize.minimize, and the boundaries for each parameter are specified in Tab. 3. We emphasize that the analytic model is crucial for us to perform the computation-heavy optimization on hundreds of datasets. Plots of the $q_{0}$ distribution at specific $M_{\mathrm{sh}}^{\mathrm{true}}$ values, as well as further discussion of the distribution shape can be found in App. A.
Figure 7: Test statistic $q_{0}$ for subhalo detectability (Eq. 26) as a function of true subhalo mass $M_{\mathrm{sh}}^{\mathrm{true}}$. Here we consider a stream similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$, $\lambda=100~\mathrm{deg}^{-1}$) for all four observational scenarios. The solid lines indicate the median value of $q_{0}$ for each truth value of the subhalo mass, $M_{\mathrm{sh}}^{\mathrm{true}}$, and the bands indicate the $\pm 1\sigma$ region. The dashed lines indicate the threshold to reject the null hypothesis of no impact at 95% CL.
@@
In order to obtain the minimum detectable subhalo mass at 95% confidence level, we use the probability distribution for $q_{0}$ when the data $\nu_{ij}^{\mathrm{dat}}$ is generated from $M_{\mathrm{sh}}^{\mathrm{true}}=0$ (no impact). The threshold for detectability is given by the 95th percentile value $q_{0}^{95\%}|_{M_{\mathrm{sh}}^{\mathrm{true}}=0}$, shown as the dashed lines in Fig. 7. For each observational scenario, the intersection of $q_{0}^{95\%}|_{M_{\mathrm{sh}}^{\mathrm{true}}=0}$ with the distribution of $q_{0}$ values for nonzero $M_{\mathrm{sh}}^{\mathrm{true}}$ gives the median and $\pm 1\sigma$ band for the minimum detectable subhalo mass at 95% CL.
Repeating this procedure for different streams, we can obtain the dependence of subhalo detectability on stream properties. Fig. 8 shows the minimum detectable subhalo mass as functions of stream width $\sigma_{\theta}$ (left), stream distance $r_{0}$ (middle) and stream density $\lambda$ (right) respectively. In each panel, we vary one stream property and keep the other two fixed. The fixed properties always take the values similar to GD-1 ($\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$ and $\lambda=$100 deg-1). The points in the plots correspond to the median while the error bars correspond to $\pm 1\sigma$. It can be seen that the minimum detectable mass increases as the stream width $\sigma_{\theta}$ increases, as the distance to the Galactic Center $r_{0}$ increases, or as the stream density $\lambda$ decreases. In addition, improvements in observational scenarios can significantly improve detectability.
@@
Figure 9: Decomposition of the test statistic $q_{0}$ into contributions from different observables. We assume a cold stream with $\sigma_{\theta}=0.2^{\circ}$, $r_{0}=10~\mathrm{kpc}$, $\lambda=100~\mathrm{deg}^{-1}$ and $q_{0}$ is evaluated on Asimov data sets, with no errors added, to show a representative breakdown. Top: We assume the minimum detectable subhalo mass in each observational scenario. Bottom: High subhalo mass $M_{\mathrm{sh}}=10^{8}~\mathrm{M}_{\odot}$ for all scenarios. In both panels, the total $q_{0}$ is normalized to 1 so that breakdowns are more comparable across different scenarios.
@@ 4.2 Fitting function for minimum detectable subhalo mass
$M_{\mathrm{sh}}^{\mathrm{min}}=\left(\frac{\sigma_{\theta}}{\mathrm{deg}}\right)^{c_{\sigma_{\theta}}}\left(\frac{r_{0}}{\mathrm{kpc}}\right)^{c_{r_{0}}}\left(\frac{\lambda}{\mathrm{deg}^{-1}}\right)^{c_{\lambda}}10^{c_{\mathrm{base}}}~\mathrm{M}_{\odot}$ (27)
@@
$c_{\sigma_{\theta}}$ $c_{r_{0}}$ $c_{\lambda}$ $c_{\mathrm{base}}$
Gaia $1.19^{\pm 4\%}$ $1.91^{\pm 3\%}$ $0.82^{\pm 4\%}$ $7.39^{\pm 1\%}$
DESI+Gaia $1.16^{\pm 4\%}$ $1.9^{\pm 3\%}$ $0.84^{\pm 4\%}$ $7.42^{\pm 1\%}$
Via+Gaia+LSST $1.07^{\pm 5\%}$ $1.33^{\pm 5\%}$ $0.83^{\pm 5\%}$ $7.31^{\pm 1\%}$
Via+Gaia+LSST10 $0.98^{\pm 9\%}$ $1.0^{\pm 12\%}$ $0.81^{\pm 8\%}$ $7.32^{\pm 2\%}$
@@
From Tab. 4, we see that the detectability depends strongly on the stream properties for all observational scenarios. For example, an approximate scaling is given by $M_{\mathrm{sh}}^{\mathrm{min}}\propto\sigma_{\theta}^{1.2}r_{0}^{1.9}\lambda^{-0.8}$ for the Gaia scenario and $M_{\mathrm{sh}}^{\mathrm{min}}\propto\sigma_{\theta}^{0.98}r_{0}^{1.0}\lambda^{-0.8}$ for LSST 10 year sensitivity. In particular, the coefficient for the stream distance $r_{0}$, $c_{r_{0}}$, has very different values for the observational scenarios. The detectability under the first two scenarios degrades significantly ($c_{r_{0}}\approx 2$) with the increase of $r_{0}$, while in the latter two scenarios, the detectability degrades much more mildly ($c_{r_{0}}\approx 1$) with the increase of $r_{0}$. This is because the number of observable stars in Gaia decreases much faster with $r_{0}$ than for LSST 10 year sensitivity (see Fig. 3). The other coefficients have little or no dependence on observational scenario.
The application of this fit to a specific observed stream may not be that accurate because of the assumptions made in this work, including a) the circular orbit for the stream, b) the default parameters for the subhalo impact, c) the assumption of fixed age and metallicity used in the MIST isochrones, and d) the usage of fixed angular bin size of $2^{\circ}$ which might not be optimal for farther away streams, where impacts typically have smaller angular size. However it gives us an effective way to quickly identify promising streams with potential for subhalo detection, from a large number of observed streams. We can always perform more accurate tests on the smaller set of promising streams that are selected from this approximate method.